# FMNV: A Dataset of Media-Published News Videos for Fake News Detection

### v2

@@ 4.1.3 Audio
We leverage semantic information from audio by first transcribing the audio stream into text using Whisper. The transcribed text is tokenized via BERT’s WordPiece tokenizer following the same pipeline as the title module, starting with a [CLS] token and ending with [SEP]. The sequence is padded/truncated to a fixed length $l_{a}=64$. Processed by the shared BERT encoder, the tokens are transformed into contextualized embeddings. We extract the final hidden states of the first 64 tokens as audio-derived text features $F_{a}=[w_{1},…,w_{l_{a}}]$, where parameter sharing with the title encoder ensures cross-modal representation alignment.
Figure 6: The main structure of Co-Attention.
@@ 4.2 Feature Aggregation
To effectively integrate information across modalities, we adopt a hierarchical co-attention mechanism inspired by Qi et al. [3], the main structure of Co-Attention is shown in figure 6. which enables deep interaction between textual, auditory, and visual signals. This design helps the model capture fine-grained cross-modal alignments and contextual dependencies.
We first apply co-attention between the title features $F_{t}\in{R}^{L_{t}\times d}$ and the audio transcription features $F_{a}\in{R}^{L_{a}\times d}$. Both modalities are projected into a shared space using linear layers. Then, each modality attends to the other by computing multi-head attention in both directions:
@@
Specifically, $F_{t}$ acts as query and attends to $F_{a}$ (as key and value), and vice versa. The outputs from both directions are integrated with residual connections and layer normalization:
$F_{t}^{\prime}$ $=\text{LayerNorm}(F_{t}+\text{Attn}_{a\rightarrow t})$ (6)
$F_{a}^{\prime}$ $=\text{LayerNorm}(F_{a}+\text{Attn}_{t\rightarrow a})$ (7)
After audio-text fusion, we use the enhanced title representation $F_{t}^{\prime}$ to interact with visual frame features $F_{v}\in{R}^{L_{v}\times d}$ using the same co-attention procedure. This produces updated features $F_{t}^{\prime\prime}$ and $F_{v}^{\prime}$:
@@
The final fused representation is obtained by concatenating the three enhanced modalities:
@@
This fused vector is linearly projected and passed into a Transformer encoder with $L$ layers to capture higher-order interactions, followed by a multi-layer perceptron (MLP) for final classification:
@@
By gradually enhancing the title features through audio and visual interactions, our model captures richer multi-modal semantics and improves its ability to detect subtle inconsistencies indicative of fake news.
@@ 5 Experiments
In this section, we conduct experiments to evaluate the performance of FMNVD and other fake news detection methods on our newly proposed FMNV dataset.
@@ 5.1 Baseline
5.1.1 Pre-trained Models
1) BERT+ResNet We adopt a dual-stream architecture where the text modality is encoded using BERT [17], a transformer-based language model, to capture contextual semantics. For visual content, ResNet-50 [18] is utilized to extract spatial features from video frames or images. The text and visual features are concatenated or fused via cross-modal transformers to predict authenticity. 2)CLIP [19], a vision-language model pre-trained on 400 million image-text pairs. We fine-tune CLIP’s text and image encoders on the FMNV and compute similarity scores between video frames and accompanying text.
5.1.2 Neural Network-based Models
@@ 5.2 Experimental Settings
The models are optimized using the Adam optimizer [20] with an initial learning rate of $1\times 10^{-3}$. To ensure a rigorous evaluation, we measure four classification metrics: Accuracy, F1-score, Precision, and Recall. All experiments are conducted on an NVIDIA GeForce RTX 2080 Ti GPU with a batch size of 128 for 30 epochs. The cross-entropy loss function is minimized during training, defined as:
$\mathcal{L}_{\text{CE}}=-\frac{1}{N}\sum_{i=1}^{N}\sum_{c=1}^{C}y_{i,c}\log(p_{i,c})$ (11)
where $N$ is the batch size, $C$ is the number of classes, $y_{i,c}$ denotes the ground-truth label, and $p_{i,c}$ represents the predicted probability for class $c$.
@@ 5.3 Experiments Results
FMNVD CD CE VS CA ALL 64.00 88.15 86.67 80.95 74.17 58.64 87.49 84.37 69.12 72.68 62.06 86.63 82.30 66.88 72.52 59.17 90.56 90.00 75.00 72.89
To validate the effectiveness of our proposed baseline and other FNVD methods on the FMNV dataset, we designed the following experiments: We categorize the fake samples in the test set into four groups - CD, CE, SV, and CA - corresponding to four distinct data augmentation methods. We evaluate several models on each categorical test subset to analyze their detection capabilities for different types of forged news videos. Additional evaluation was conduct on the entire test set to verify the overall effectiveness of our proposed method. The experimental results shown in Table 3 lead us to the following conclusions: 1) All three neural network-based methods achieved accuracies above 0.72 on the FMNV dataset, with FMNVD demonstrating superior performance among the five baseline models by attaining 0.7417, which confirms both the generalization capability of FMNV and the competitive advantage of FMNVD. 2) The evaluated methods exhibited excellent detection performance for fake news videos containing Cherry-picked Editing and Synthetic Voiceover manipulations. However, they show relatively poor detection accuracy for Contextual Dishonesty-type videos. This limitation primarily stems from the fact that Contextual Dishonesty samples are generated through subtle phrase-level modifications in video titles, which creates semantically nuanced variations that current models struggle to detect effectively. This presents a significant challenge for future methodological improvements.
Table 4: Ablation study on FMNVD. (a) Modality removal. (b) Hyperparameter sensitivity.
(a) Modality Ablation
@@
All 74.17 72.68 72.52 72.89
w/o Title 65.42 65.37 68.16 68.78
w/o Frames 71.25 70.47 70.43 71.67
w/o Audio 63.75 58.83 60.22 58.78
(b) Hyperparameter Sensitivity
Setting Accuracy F1 Precision Recall
Patch size = 16 73.84 72.23 72.00 72.50
Seq. len. = 32 74.17 71.91 72.46 71.56
Seq. len. = 64 73.75 72.96 72.80 74.11
$\tau=0.3$ 74.17 71.12 72.82 70.44
$\tau=0.7$ 72.50 70.67 70.67 70.67
5.4 Ablation Study
To assess both the contribution of each modality and the robustness of our framework to key hyperparameters, we perform two sets of ablations on FMNVD:
(a) Modality Ablation We individually remove each modality (title, frames, audio) and its associated co-attention layers, and measure the resulting drop in performance. As shown in the top block of Table4, all three modalities are vital, with audio having the largest impact, followed by title and then visual frames. This phenomenon primarily stems from our implementation strategy where audio signals are first converted into text through Whisper for processing. The transcribed audio text inherently contains more linguistic content than video titles, and such extended textual information proves more impactful than shorter textual inputs (titles) and visual features in our framework.
(b) Hyperparameter Sensitivity To verify the stability of our detection framework, we further vary three core hyperparameters:
  • CLIP patch size (e.g. 16)
  • Maximum sequence length for text/audio features (e.g. 32 vs. 64 tokens)
  • Confidence threshold $\tau$ for filtering low-confidence predictions (e.g. 0.3 vs. 0.7)
The bottom block of Table 4 provides the performance under each setting, demonstrating that our model maintains stable accuracy and F1 across a reasonable range of each hyperparameter.
@@ 6 Conclusion
In this paper, we conduct an empirical analysis for fake news video categorization and construct a novel fake news video dataset (FMNV) using large model-based data augmentation. We provide a comprehensive description of FMNV’s creation process with in-depth investigations into data entity attributes and distribution characteristics. To assess the dataset’s validity, we introduce a new baseline model FMNVD and perform numerical experiments to evaluate various baseline performances, thereby demonstrating FMNV’s generalization capability. Future work will explore additional promising approaches to enhance the dataset’s predictive power, with the expectation that FMNV will pave the way for advanced developments in MFND research.

### v3

@@ 3.1.3 Audio
Audio is transcribed using Whisper and tokenized via BERT, following the title module. Sequences are padded/truncated to $l_{a}=64$, then encoded using the same BERT model. The first 64 token embeddings form $F_{a}=[w_{1},\dots,w_{l_{a}}]$. Parameter sharing ensures alignment between title and audio representations.
@@ 3.2 Feature Aggregation
To integrate information across modalities, we adopt a hierarchical co-attention mechanism inspired by Qi et al. [3], which enables deep interactions among textual, auditory, and visual features.
We first apply bi-directional co-attention between title features $F_{t}\in{R}^{L_{t}\times d}$ and audio transcription features $F_{a}\in{R}^{L_{a}\times d}$, both projected into a shared space. Multi-head attention is computed in both directions:
@@
with residual connections and layer normalization:
$F_{t}^{\prime}$ $=\text{LayerNorm}(F_{t}+\text{Attn}{a\rightarrow t})$ (6)
$F_{a}^{\prime}$ $=\text{LayerNorm}(F_{a}+\text{Attn}{t\rightarrow a})$ (7)
Next, the enhanced title features $F_{t}^{\prime}$ attend to visual frame features $F_{v}\in{R}^{L_{v}\times d}$ using the same co-attention strategy, producing updated representations $F_{t}^{\prime\prime}$ and $F_{v}^{\prime}$:
@@
The final fused vector is formed by concatenating the three enhanced modalities:
@@
and passed through a Transformer encoder followed by an MLP for classification:
@@
This hierarchical fusion enables the model to capture fine-grained cross-modal semantics and improves its robustness to subtle manipulations in fake news content.
@@ 4.2 Experimental Settings
The models are optimized using the Adam optimizer [16] with an initial learning rate of $1\times 10^{-3}$. To ensure a rigorous evaluation, we measure four classification metrics: Accuracy, F1-score, Precision, and Recall. All experiments are conducted on an NVIDIA GeForce RTX 2080 Ti GPU with a batch size of 128 for 30 epochs. The cross-entropy loss function is minimized during training.
@@ 4.3 Experiments Results
FMNVD CD CE VS CA ALL 64.67 87.41 86.67 77.14 74.17 60.05 86.75 84.60 61.44 73.76 62.82 86.00 82.61 60.34 74.17 60.28 90.00 91.11 64.44 75.78
To validate the effectiveness of our proposed method and other baselines on the FMNV dataset, we conducted experiments by dividing the fake samples in the test set into four categories: Contextual Dishonesty (CD), Cherry-picked Editing (CE), Synthetic Voiceover (SV), and Cross-modal Attribution (CA), corresponding to four distinct manipulation strategies. Each model was evaluated on these subsets to analyze their detection performance across manipulation types. Additionally, we assessed performance on the full test set to examine overall effectiveness.
As summarized in Table 3, all three models achieved over 0.72 accuracy, with our proposed FMNVD model attaining the highest score of 0.7417. This confirms both the generalization capability of the FMNV dataset and the competitive performance of FMNVD. Notably, the models perform well on CE and SV manipulations, while detection on CD samples remains challenging due to their subtle, phrase-level semantic alterations. This highlights a key limitation in current methods and points to future directions for improving fine-grained semantic reasoning.
4.4 Ablation Study
We further conduct a multimodal ablation study using FMNVD on the dataset to analyze the contributions of different modalities in FMNV. Specifically, we individually remove the title, video, and audio modalities from the data while simultaneously eliminating their corresponding cross-modal attention layers in the model. The experimental results presented in Table 4 demonstrate that all three modality ablations caused notable performance degradation in the baseline model. The relative importance of different modalities to model performance can be ranked as audio >title >frames. This phenomenon primarily stems from our implementation strategy where audio signals are first converted into text through Whisper for processing. The transcribed audio text inherently contains more linguistic content than video titles, and such extended textual information proves more impactful than shorter textual inputs (titles) and visual features in our framework.
Table 4: Results of ablation study.
@@ 4.4 Ablation Study
All 74.17 73.76 74.17 75.78
w/o Title w/o Frames w/o Audio 65.00 72.50 60.00 64.00 71.94 57.87 64.06 72.11 57.81 64.89 73.56 58.00
@@ 5 Conclusion
In this paper, we conduct an empirical analysis for fake news video categorization and construct a novel fake news video dataset (FMNV) using large model-based data augmentation. We provide a comprehensive description of FMNV’s creation process with in-depth investigations into data entity attributes and distribution characteris-tics. To assess the dataset’s validity, we introduce a new baseline model FMNVD and perform numerical experiments to evaluate various baseline performances, thereby demonstrating FMNV’s generalization capability. Future work will explore additional promising approaches to enhance the dataset’s predictive power, with the expectation that FMNV will pave the way for advanced developments in MFND research.